OpenStudy (anonymous):

Given two independent events X and Y, with P(XlY)=0.45 and the P(X and y)=0.25. what is the P(Y)?

OpenStudy (anonymous):

\[P(X/Y) = P(X and Y) / P(Y)\] \[\rightarrow P(Y) = P(X and Y) / P(X/Y) \] = 0.25 / 0.45 = 0.55